OpenStudy (anonymous):

Solve the following for x: 2x over x+1=1/3-2 over x+1 ( Solve for x, than check your answer is it extraneous)?

OpenStudy (anonymous):

x=-1,-5/6

OpenStudy (anonymous):

Thanks, BVEL.:)

OpenStudy (agreene):

\[\frac{2x}{x+1}=\frac{\frac{1}{3-2}}{x+1}\] is that what you are asking? if it is your answer is 1/2

OpenStudy (anonymous):

No, the second equation is 1/3 - 2/x+1 the first equation is correct.

OpenStudy (agreene):

\[\frac{2x}{x+1}=\frac{1}{3}-\frac{2}{x+1}\] This has no solutions.

OpenStudy (anonymous):

Yes, I am aware it has no solutions but it still asks to figure out x?